AMA INTERNATIONAL UNIVERSITY-BAHRAIN Salmabad, Kin
Answer #1

150V c.

Using KVL in loop ABCDA , we get the equation as :

(20 ohm) I + (-2Vo) + (25 ohm) I + (-Vo) + (-150) =0

20I - 2Vo + 25I - Vo - 150 =0              Eq-1

In branch CD of the circuit , Voltage = Vo , current = I and resistance = 20

so using ohm's law , Vo = 20I

replacing ''Vo'' by ''20I'' in Eq-1 ,

20I - 2Vo + 25I - Vo - 150 =0      

20I - 2(20I) + 25I - 20I - 150 =0

20I - 40I + 25I - 20I -150 =0

-15I -150 =

-15I = 150

I = -10 A

so Vo = 20 I = 20 (-10) = -200 Volts             (since I = -10 A)

B)

circuit1      Зол point Q and S cań b joined 10 ohm and 50 ohm are in series sinde both are joined only at one 1on S0 SL 1 o -( DA个 ohm are

series Combination of resistance = R1 + R2

so Combination of 10 ohm and 50 ohm = 10 + 50 = 60 ohm,

similarly , combination of 70 oh and 80 ohm = 70 + 80 = 150 ohm

Diagram Now looks as

フ 3L 30 ohm and 20 ohm are in parallel since they are connected to each other at both end 51 15o1

parallel Combination of reistances = \frac{R_{1}\times R_{2}}{R_{1} +R_{2}}     ,   

So parallel combination of 30 ohm and 20 ohm = \frac{30\times 20}{30 +20} = 12 ohm

parallel combination of 40 ohm and 60 ohm = \frac{40\times 60}{40 +60} = 24 ohm

parallel combination of 60 ohm and 150 ohm = \frac{60\times 150}{60 +150} = 42.86 ohm

so Diagram now looks as

all resistances are now in series.

so total resistance is simply the sum of all.

so Rtotal = 8 ohm + 12 ohm + 24ohm + 6 ohm + 42.86 ohm + 4 ohm = 96.86 ohm